Grade 12Algebra

the radius of circle toching parabola y2=x at (1,1) and having directrix of y2=x as its normal is?

Nishtha Gahlot
9 Years agoGrade 12

1 Answer

Allu Sidhu
9 Years ago
The given parabola is in standard form..With a=1/4..The equation of directrix is x+a=0..X+1/4=0We know that normal to a circle passes through center of circle ..Therefore question is reduced to finding perpendicular distance of (1,1) to the line that is 5/4Therefore radius is 5/4